Aniline, C6H5NH2 is a weak base. A 0.3 M aniline solution was prepared and it was determined that the aniline had undergone 0.5% ionisation. Calculate the pH of this solution.
1
Expert's answer
2017-07-10T05:32:15-0400
5% = 0.05
pH = 14 - pOH
pOH = -lg(OH-) = -lg(0.3*0.05) = -lg(0.015) = 1.82
pH = 14 - 1.82 = 12.18